OpenStudy (anonymous):

simplify the expression completely... (t/(t+3)) + (4t/(t-3)) - (18/(t^2 - 9)) please explain too...

OpenStudy (anonymous):

i arrived to (5t-6)/(t-3) can i take it any further?

OpenStudy (gw2011):

You need to find a common denominator. Also, (t^2-9)=(t+3)(t-3). Now you get the following: [t/(t+3)] + [4t/(t-3)] - [18/(t+3)(t-3)] The common denominator is: (t+3)(t-3) [(t)(t-3)+(4t)(t+3)-18]/(t+3)(t-3) (5t^2+9t-18)/(t+3)(t-3) (5t-6)(t+3)/(t+3)(t-3)=(5t-6)/(t-3) Your answer is correct.

OpenStudy (anonymous):

thank you!!

OpenStudy (gw2011):

You're welcome